### v1

@@ 3.0.2 Implementation
All image encoders are 3D ResNet-50 models implemented using the MONAI framework [4]. The embedding dimensions are set to 2048 for the encoder and 128 for the projection head, following SimCLR [5]. The similarity threshold $h$ is set to 0.05, the regularization parameter $\lambda$ to 0.5, and the temperature parameter $\tau$ to 0.1. The representations are pre-trained for 10 epochs using a batch size of 512 and a learning rate of 1e-3. For downstream tasks, the projection head is removed, and a linear layer is added on top of the image encoder, with an output size corresponding to the number of classes. All downstream tasks are fine-tuned for 35 epochs. For pre-training of MMCL [8] and SimCLR, we follow the original settings described by the authors, with modifications to match the 10-epoch training schedule used by our method.
@@ 4 Results
4.1 Representations Strength Comparison
@@
Table 2: Zero-shot demographics prediction. Sex is measured using accuracy, while the other attributes using MAE. The * corresponds to attributes that were not included at pre-training.
Sex BMI* Height* Weight*
@@
SimCLR [5] 75.95 2.56 6.2 8.54
MMCL [8] 92.71 2.98 5.12 8.82
TGV (Ours) 98.34 2.31 4.89 7.41
Figure 2: Performance of the models under low data regime.
We compare the performance of our method, with and without image augmentations, against a supervised ResNet50 [9] and two self-supervised representation learning methods, image-only SimCLR [5] and image-tabular MMCL [8], across two downstream task categories. First, we report the performance on multi-label CAD prediction using the area under the receiver operating characteristic curve (AUC), as it provides a robust evaluation metric for datasets with a low frequency of non-healthy samples. Second, we evaluate cardiac phenotype prediction using mean absolute error (MAE) across six attributes: left ventricular ejection fraction (LVEF), left ventricular end-diastolic mass (LVEDM), left ventricular end-diastolic volume (LVEDV), right ventricular ejection fraction (RVEF), right ventricular end-diastolic volume (RVEDV), and myocardial end-systolic volume (MYOESV). These attributes are selected to comprehensively assess models’ performance across two characteristic phases of the cardiac cycle and to ensure coverage of the three main anatomical regions of the heart. Table 1 reports the results for both zero-shot prediction and fine-tuning.
TGV without image augmentations (denoted ours in the table) achieves the best performance across all the tasks on both fine-tuning and zero-shot predictions, demonstrating the strength of tabular data guidance in training medical visual representations. This shows that defining clinically meaningful pairs yields stronger representations than image augmentations-based pair sampling. Our comparison of TGV with and without image augmentations shows that not using image augmentations improves the strength of the representation. Further, tabular guidance outperforms tabular supervision, suggesting that training a joint embedding could limit the encoder’s ability to learn directly from the images themselves, due to a strong supervisory signal being provided by the other modality. In contrast, tabular guidance directs the model to learn clinically meaningful information embedded in the tabular attributes directly from the images, resulting in a stronger visual encoder.
4.1.1 Performance under Low-Data Regimes
Table 3: Impact of parameter $\lambda$ (Eq. 1) on the transferability of the method using fine-tuning (FT) and zero-shot (ZS) predictions. $\lambda>0.5$ increases the importance of continuous attributes in the similarity calculations, while $\lambda<0.5$ puts more weight on the categorical attributes.
Training $\quad\lambda\quad$ CAD LVEF LVEDM LVEDV RVEF RVEDV MYOESV
   1 62.0 4.36 7.68 12.94 4.24 11.47 8.17
   0.75 64.39 4.36 8.82 14.22 4.4 12.16 9.35
ZS    0.5 67.97 4.08 7.65 13.7 3.98 12.38 8.16
   0.25 65.32 4.52 12.06 19.95 4.57 18.76 12.29
   0 65.44 4.55 11.18 21.46 4.6 20.54 11.4
   1 74.05 3.23 5.17 9.57 3.07 7.55 5.39
   0.75 73.83 3.32 5.47 9.74 3.16 7.65 5.59
FT    0.5 76.1 3.18 4.86 9.23 2.95 7.39 5.2
   0.25 71.39 3.39 5.89 10.28 3.3 7.92 6.19
   0 71.38 3.49 5.98 10.62 3.37 8.11 6.11
Medical data annotation can be costly and time-consuming. To evaluate the efficiency of the models in low-data regime, we sample 1% (64 samples) and 10% (642 samples) of the original classification training set and use it for fine-tuning. Fig. 2 presents the results of the models fine-tuned with limited data, including a comparison against their performance in a zero-shot setting. TGV achieves the best performance across all data regimes. Moreover, the zero-shot performance of TGV is better than or comparable to the low-data fine-tuning performance of SimCLR [5] and MMCL [8], highlighting the strength of the representations learned by our method.
4.2 Zero-Shot Extraction of Demographic Information
@@
Clinicians commonly use demographic information in their decision making. Thus, it is crucial for an image encoder to capture that information. We perform a zero-shot prediction of four demographic attributes, including sex, weight, height, and body-mass index (BMI). Sex prediction is performed as classification and measured using accuracy, while the other metrics are predicted as a value and evaluated using MAE. Only the sex attribute was included during training, meaning no information about weight, height, or BMI was incorporated to guide the training process. As shown in Table 2, TGV achieved the best performance at capturing the demographics information in its representation space. This suggests that tabular-data guidance allows the encoder to capture information encoded in the image, such as weight or height, even without explicitly incorporating those attributes at training. MMCL [8] outperforms SimCLR [5] in sex and height prediction, while SimCLR achieves better results on BMI and weight. This result further solidifies our claim that tabular data supervision might limit the encoder’s ability to capture information directly from the images by providing a strong supervisory signal in a form of a different modality, reducing the strength of the image encoder itself.
4.3 Impact of Attribute Weighting on the Embedding Transferability
@@
We evaluate the impact of weighting tabular guidance towards categorical or continuous data by experimenting with different values of the hyperparameter $\lambda$ from Eq. 1. The results are presented in Table 3. The model trained with $\lambda=0.5$ achieved the best results, outperforming the other models in nearly all zero-shot and all fine-tuning tasks, suggesting that the best performance is achieved when categorical and continuous attributes contribute to the guidance in the same amount. Strong influence of categorical attributes ($\lambda<0.5$) yields the poorest results. As our dataset has a very small frequency of samples with diseases (8%) and the categorical attributes consist mostly of those describing presence of CAD, there might not be enough variability between the attributes to properly guide the training. Increased influence of continuous attributes ($\lambda>0.5$) lowers the performance in comparison to equal guidance, but not as much as strong categorical guidance. This suggests that while continuous attributes retain variability across samples, their differences may not provide sufficiently descriptive signals for tabular guidance.
@@ 5 Conclusion
We present TGV, a contrastive learning paradigm leveraging tabular data to generate clinically meaningful pairs for training of visual representations. Our approach outperforms augmentations-based image only contrastive learning and tabular-supervision on CAD classification and cardiac phenotype prediction, highlighting the strength of our approach in a medical setting. Additionally, we propose a zero-shot prediction method compatible with unimodal image representations, overcoming a crucial limitation of those representations. Through zero-shot experiments, we show that our method outperforms previous approaches at capturing demographic information in its representation space. TGV can be leveraged to train medical foundation models grounded on rich clinical information, paving the way for more robust and generalizable medical models.

### v2

@@ 4.2 Implementation
All image encoders are 3D ResNet50 models implemented using the MONAI framework [5]. The embedding dimensions are set to 2048 for the encoder and 128 for the projection head, following SimCLR [6]. The similarity threshold $h$ is set to 0.05, the regularization parameter $\lambda$ to 0.5, and the temperature parameter $\tau$ to 0.1. The models are all pretrained for 10 epochs to ensure fair comparison. We use batch size of 512 and learning rate of 1e-3. For downstream tasks, the projection head is removed, and a linear layer is added on top of the image encoder, with an output size corresponding to the number of classes. All downstream tasks are fine-tuned for 35 epochs. Baseline pretraining follows the authors’ original settings [18, 6, 16, 40, 7], modified to match our 10-epoch schedule. See SM Section 2 for details. The supervised models are trained on the training sets prepared for the fine-tuning.
@@ 5 Results
Figure 3: Performance of the models under different amount of training data on LVEF prediction (left, lower is better) and multilabel CAD classification (right, higher is better) tasks. The performance is evaluated using fine-tuning (solid lines), linear probing (dashed lines), and zero-shot prediction (stars).
5.1 Tabular Guidance Outperforms Image Augmentation & Tabular Supervision
We compare the performance of our method against a supervised ResNet50 [20] and five contrastive learning baselines: four image-only methods, SimCLR [6], SimSiam [8], Barlow Twins [40], and BYOL [16], and one image-tabular method, MMCL [18], which uses tabular supervision (as illustrated in Fig. 2). All methods are evaluated across two downstream task categories. First, we report the performance on multilabel CAD classification using the area under the receiver operating characteristic curve (AUC), as it provides a robust evaluation metric for datasets with a low frequency of non-healthy samples. Second, we evaluate cardiac phenotype prediction using mean absolute error (MAE) across six attributes: left ventricular ejection fraction (LVEF), left ventricular end-diastolic mass (LVEDM), left ventricular end-diastolic volume (LVEDV), right ventricular ejection fraction (RVEF), right ventricular end-diastolic volume (RVEDV), and myocardial end-systolic volume (MYOESV). These attributes are selected to comprehensively assess the models’ performance across two characteristic phases of the cardiac cycle and to ensure coverage of the three main anatomical regions of the heart. Table 1 reports the results for both zero-shot prediction and fine-tuning. The zero-shot results are reported as a mean over the three representative sets $P$. To evaluate the robustness of the zero-shot predictions, we report the standard deviations over the three sets in SM Table 4, while SM Table 3 reports results over different sizes of the three sets.
TGV achieves the best performance across all the tasks on both fine-tuning and zero-shot predictions, demonstrating the strength of tabular data guidance in training visual representations. This shows that defining clinically meaningful pairs yields stronger representations than image augmentation-based pair sampling, further showing the importance of incorporating the semantic information into the pair sampling process. Additionally, tabular guidance also outperforms tabular supervision, suggesting that training a joint embedding could limit the encoder’s ability to learn directly from the images themselves due to a strong supervisory signal being provided by the other modality. In contrast, tabular guidance directs the model to learn clinically meaningful information embedded in the tabular attributes directly from the images, resulting in a stronger visual encoder. Furthermore, to show that TGV can generalize to other datasets and domains, we provide experiments on the DVM [23] dataset in the SM Section 4.
5.2 TGV is Effective under Limited Data
Data annotation can be costly and time-consuming, especially in the medical domain. To evaluate the efficiency of the models in low-data regime, we sample 1% (CAD: 64, LVEF: 50 samples) and 10% (CAD: 642, LVEF: 500 samples) of the original classification and phenotype prediction training sets and use it for fine-tuning and linear probing. The validation and test sets are unchanged. Fig. 3 presents the results of the models trained with limited data. For clarity, we only include MMCL [18] and SimCLR [6], while the results of the other baselines are presented in the SM Fig. 1. TGV achieves the best performance across all data regimes for both CAD classification and LVEF prediction. At the 10% data regime, TGV matches or outperforms SimCLR and MMCL trained on the full dataset for CAD classification, under both fine-tuning and linear probing. These findings highlight the strength of the representation learned by our method, demonstrating its utility in settings with limited data, such as rare diseases.
5.3 k-NN Allows Unimodal Zero-Shot Prediction
The lack of zero-shot abilities is one of the crucial limitations of unimodal representations. We evaluate the performance of the unimodal zero-shot approach in comparison with fine-tuning and linear probing. While fine-tuning on the full dataset achieves the highest overall performance, our zero-shot method often surpasses fine-tuning when labeled data is limited, and consistently outperforms linear probing across all data regimes for LVEF prediction (Fig. 3). This strong LVEF prediction performance is likely because LVEF is a continuous and widely available measure, enabling well-clustered representations (Fig. 4) that support accurate zero-shot regression. Although CAD classification is more challenging due to class imbalance (8% positive cases), zero-shot performance stands competitive on this task in low-data regimes.
Figure 4: t-SNE visualization of the feature embedding generated with TGV. Sex, LVEF, and LVEDV have been included as attributes for calculation of tabular similarity during training. Height was not included.
5.4 TGV Builds Clinically Aware Embeddings
Does TGV capture clinically meaningful structure in its latent space? To assess the quality of the learned representations, we visualize the t-SNE of test set embeddings, colored by sex, height, LVEF, and LVEDV. The results for TGV are shown in Fig. 4 (see SM Fig. 3 for SimCLR [6] and MMCL [18]). TGV is able to generate representations where similar phenotypes are clustered at close locations. Furthermore, height is not included in the tabular similarity calculations at training, yet we can still see proper clustering of these attribute. The generated embeddings exhibit a primary separation based on sex, which is consistent with established cardiac physiological differences between sexes [32]. These observations suggest that TGV captures underlying phenotype structure in a clinically meaningful way, even beyond explicitly provided training signals.
@@ 5.4 TGV Builds Clinically Aware Embeddings
Table 2: Zero-shot demographics prediction. Sex is measured using accuracy, while the other attributes using MAE. The * corresponds to attributes that were not included at pretraining.
Sex $\uparrow$ BMI* $\downarrow$ Height* $\downarrow$ Weight* $\downarrow$
@@
SimCLR [6] 75.04 2.54 6.24 8.43
BYOL [16] 69.72 3.33 7.02 10.44
SimSiam [8] 68.19 2.86 6.93 8.80
Barlow Twins [40] 72.90 3.16 7.02 10.88
MMCL [18] 92.60 2.98 5.12 8.82
TGV (Ours) 98.16 2.39 4.88 7.44
@@
Can tabular guidance learn demographics? Clinicians commonly use demographic information in their decision making. Thus, it is crucial for an image encoder to capture that information. We perform a zero-shot prediction of four demographic attributes, including sex, weight, height, and body-mass index (BMI). Sex prediction is performed as classification and measured using accuracy, while the other metrics are predicted as a value and evaluated using MAE. Only the sex attribute was included during training, meaning no information about weight, height, or BMI was incorporated to guide the training process. As shown in Table 9, TGV achieved the best performance at capturing the demographics information in its representation space. This suggests that tabular-data guidance allows the encoder to capture information encoded in the image, such as weight or height, even without explicitly incorporating those attributes at training. MMCL [18] outperforms image augmentation methods in sex and height prediction, but fails to image augmentation-based SimCLR [6] on BMI and weight. This result further solidifies our claim that tabular data supervision might limit the encoder’s ability to capture information directly from the images by providing a strong supervisory signal in a form of a different modality, reducing the strength of the image encoder itself.
5.5 Ablation Studies
Table 3: Performance under different values of threshold h (Eq. 6), which controls the multi-pair approach.
Threshold h CAD $\uparrow$ LVEF $\downarrow$
ZS FT ZS FT
0.0 56.35 72.06 4.90 3.89
0.05 68.70 76.1 4.08 3.18
0.1 67.17 74.41 4.37 3.47
0.2 67.60 74.63 4.38 3.33
Table 4: Impact of parameter $\lambda$ (Eq. 3) on the transferability of TGV for downstream tasks using fine-tuning (FT) and zero-shot (ZS) predictions. $\lambda>0.5$ emphasizes continuous attributes (Con), while $\lambda<0.5$ favors categorical fields (Cat) in similarity computations.
CAD $\uparrow$ LVEF $\downarrow$ LVEDM $\downarrow$ LVEDV $\downarrow$ RVEF $\downarrow$ RVEDV $\downarrow$ MYOESV $\downarrow$
$\lambda$ ZS FT ZS FT ZS FT ZS FT ZS FT ZS FT ZS FT
Con 1 62.0 74.05 4.36 3.23 7.68 5.17 12.94 9.57 4.24 3.07 11.47 7.55 8.17 5.39
0.75 64.39 73.83 4.36 3.32 8.82 5.47 14.22 9.74 4.4 3.16 12.16 7.65 9.35 5.59
0.5 67.97 76.1 4.08 3.18 7.65 4.86 13.7 9.23 3.98 2.95 12.38 7.39 8.16 5.2
0.25 65.32 71.39 4.52 3.39 12.06 5.89 19.95 10.28 4.57 3.3 18.76 7.92 12.29 6.19
Cat 0 65.44 71.38 4.55 3.49 11.18 5.98 21.46 10.62 4.6 3.37 20.54 8.11 11.4 6.11
One-to-one or many-to-many correspondences in contrastive learning? Assuming a perfect one-to-one correspondence often hinders the performance of contrastive learning. We evaluate the effectiveness of our multi-pair approach introduced via a hyperparameter $h$ (Eq. 6) by performing experiments on the effect of different threshold values on the downstream performance of CAD classification and LVEF prediction. At $h=0$ each anchor is only paired with one instance, resulting in a single pair approach. Increasing the value of $h$ results in more positive pairs per anchor. As presented in Table 3, TGV achieves the best performance at $h=0.05$. The single-pair approach ($h=0$), achieves the worst performance, showing the need for many-to-many approach, rather than a simple one-to-one correspondence. However, setting the value of $h$ too high could result in too many pairings, causing the model to represent overly broad or noisy relationships between samples. This can lead to less discriminative embeddings, as the model may struggle to maintain sufficient uniformity in the representation space and risk collapsing semantic distinctions between different samples. In extreme cases, this could result in a form of representation collapse, where embeddings for diverse samples become undesirably similar, undermining the effectiveness of the learned representations. On the other hand, setting the parameter too low could miss on semantic similarity between the samples and result in false negatives. The value of $h$ should be tuned based on a given dataset and its underlying relationships between samples.
Which tabular attributes contribute most to the embedding quality? To evaluate which features are of most importance for tabular guidance, we evaluate the impact of weighting tabular guidance towards categorical or continuous data by experimenting with different values of the hyperparameter $\lambda$ from Eq. 3. The results are presented in Table 4. Values of $\lambda$ below 0.5 increase the influence of categorical features, while $\lambda$ above 0.5 puts stronger emphasis on continuous data. Continuous attributes mostly consist of cardiac phenotype features and describe the function and structure of the heart. Many of them could be described as morphological information. Categorical data mostly describes the presence of cardiovascular disorders and demographic information.
@@ 5.5 Ablation Studies
The model trained with $\lambda=0.5$ achieved the best results, outperforming the other models in nearly all zero-shot and all fine-tuning tasks, suggesting that the best performance is achieved when categorical and continuous attributes contribute to the guidance in the same amount. Strong influence of categorical attributes ($\lambda<0.5$) yields the poorest results. As our dataset has a very small frequency of samples with diseases (8%) and the categorical attributes consist mostly of those describing the presence of CAD, there might not be enough variability between the samples to properly guide the training. Increased influence of continuous attributes ($\lambda>0.5$) lowers the performance in comparison to equal guidance, but not as much as strong categorical guidance. This suggests that while continuous attributes retain variability across samples, their differences may not provide sufficiently descriptive signals for tabular guidance. Furthermore, a better performance under strong continuous guidance is consistent with Hager at al. [18], where morphological features improved embeddings. The parameter $\lambda$ should be adjusted per dataset, taking into consideration the distribution of its tabular attributes.
6 Limitations
Our work proposes adapting k-NN for unimodal zero-shot prediction. However, unlike image-text zero-shot methods, our approach requires access to a representative sample of images for the target task. Acquiring such a sample may be particularly challenging in the context of rare diseases, where data availability is limited. Future work could address this limitation by generating synthetic samples of rare diseases using generative models.
Another important limitation is the demographic imbalance in the UK Biobank dataset, which predominantly consists of white subjects. Individuals from other ethnic backgrounds comprise only around 5% of the dataset. This imbalance may lead to biased predictions for underrepresented groups, particularly given known disparities in cardiac disease risk across different ethnicities [26]. Ensuring fair model performance across diverse populations remains a critical area for future research.
@@ 7 Conclusion
We present TGV: Tables Guide Vision, a contrastive learning paradigm leveraging tabular data to generate clinically meaningful pairs for training of visual representations. Our approach outperforms image augmentation-based contrastive learning and tabular supervision on coronary artery disease classification and cardiac phenotype prediction, highlighting the strength of our approach in a medical setting. Additionally, we adapt k-NN to serve as a robust zero-shot prediction approach in unimodal image representations, overcoming a crucial limitation of this setting. Through zero-shot experiments, we show that our method outperforms previous approaches at capturing demographic information in its representation space. TGV can be leveraged to train medical foundation models grounded on rich clinical information, paving the way for more robust and generalizable medical models. While we designed the method with a clinical setting in mind, experiments on natural images show that the approach is generic and can generalize to image-tabular datasets in other domains without incorporating any changes.